Question 5 20 pts WACC = 8% Period Project Project 1 -$200 -$75 $100 $100 $75 $25 $80 O Based on the information in the tableplease answer
Answer #1

NPV = CF0 + CF1/(1+R) + CF2/(1+R)2 + CF3/(1+R3)

= -200 +100/(1+8%) + 75/(1+8%)2 + 80/(1+8%)3

= -200 + 92.6 + 64.3 + 63.5 = 20.4

Ans) 20.4